OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? -5/3 5/3 -3/5 3/5

OpenStudy (zzr0ck3r):

(0-3)/(-3-2) = ?

OpenStudy (anonymous):

(-3)/(5) = 8

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1}\] Follow the gradient formula

OpenStudy (anonymous):

-3-2 = -5 not 5.

OpenStudy (anonymous):

so it 1

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1} = \frac{0-3}{-3-2}= \frac{-3}{-5}= \frac{3}{5}\]

OpenStudy (anonymous):

now which option is it??

OpenStudy (anonymous):

it would be